OpenStudy (anonymous):

If f(x)=x/4 -3 and g(x) = 4x^2 + 2x -4 find (f+g) (x) the answers are 4x^2 + 7/4x -7 4x^2 + 9/4x + 1 4x^2 + 9/4x -7 x^2 -12

OpenStudy (campbell_st):

f(x) + g(x) = (f + g)(x) so you have \[(f + g)(x) = 4x^2 + 2x + \frac{x}{4} -4 - 3\] just collect the like terms... to simplify

OpenStudy (anonymous):

^ I agree with @campbell_st

OpenStudy (anonymous):

Ok thanks!!